500gallon heating tank at 20% ethanol. You need it to be a 40% ethanol solution. You can only add 100% ethanol. How the 'blank'
Answer by stanbon(75887)

500gallon heating tank at 20% ethanol. You need it to be a 40% ethanol solution. You can only add 100% ethanol.----
---
Equation:
eth + eth = eth
0.20*500 + 1.00x = 0.40(500+x)
---
Multiply thru by 100 to get:
20*500 + 100x = 40*500 + 40x
60x = 20*500
x = (1/3)500
x = 166 2/3 gallong (amt. of 100% ethonol to be added)
=============
Cheers,
Stan H.
===========